OpenStudy (john123):

What is the vertex of the graph of the following quadratic equation -4(x+7)^2 -5,

OpenStudy (anonymous):

vertex=-b/2a -4(x+7)^2 -5 y=-4(x+7)(x+7)-5 y= -4x^2-56x-201 x=-(-56)/2(-4) x=7 y=-4(7+7)^2-5 y=-789 (7,-789)